Find how many quarts of 6% butterfat milk and 2% butterfat milk should be mixed to yield 40 quarts of 3% butterfat milk.
The mixture should contain __ quarts of the 6% butterfat milk
The mixture should contain___ quarts of the 2% butterfat mik

Answer by stanbon(75887)

Find how many quarts of 6% butterfat milk and 2% butterfat milk should be mixed to yield 40 quarts of 3% butterfat milk.
----
Equation:
fat + fat = fat
0.02x + 0.06(40-x) = 0.03*40
-----
Multiply thru by 100 to get:
2x + 6*40 - 6x = 3*40
-4x = -3*40
x = 30 quarts (amt. of 2% solution needed)
------
40-x = 10 quarts (amt. of 6% solution needed)
============================
Cheers,
Stan H.
===========



The mixture should contain __ quarts of the 6% butterfat milk
The mixture should contain___ quarts of the 2% butterfat mik